the length of a rectangle is 7 centimeters less than 4 times its width. its area is 36 square centimeters. find the dimensions of the rectangle. use the formula area= length*width
Answer by checkley79(3341)

A=LW
L=4W-7
36=(4W-7)W
36=4W^2-7W
4W^2-7W-36=0
(4W+9)(W-4)=0
W-4=0
W=4 ANS. FOR THE WIDTH.
PROOF:
36=(4*4-7)4
36=(16-7)4
36=9*4
36=36